OpenStudy (anonymous):

Find the centre and the radius of circle X^2 + Y^2 - 4X+6Y =3

OpenStudy (anonymous):

Complete the square

OpenStudy (anonymous):

\[x ^{2}+y ^{2}-4x+6y=3\]

OpenStudy (anonymous):

Okay, so if you have \(y^2+6y\), then how do you complete the square?

OpenStudy (anonymous):

\[ (y+c)^2 = y^2+2cy +c^2 \]

OpenStudy (anonymous):

if \(2c=6\) then \(c = 3\) and \(c^2=9\)

OpenStudy (anonymous):

This means that\[ (y+3)^2 = y^2+6y+9 \]

OpenStudy (anonymous):

Or in other words: \[ y^2+6y = (y+3)^2-9 \]This is how you complete the square.

OpenStudy (anonymous):

\[ x ^{2}+y ^{2}-4x+6y=3 \]Is the same as \[ x ^{2}-4x+(y+3)^2-9=3 \]

OpenStudy (anonymous):

Now you need to complete the square for \(x^2-4x\).

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

i got it thank you very much, it's been a while since i touched any question with circles, needed to brush up.

OpenStudy (anonymous):

r=4, C(2, -3)